From what height should a piece of ice fall show that it completely melt only one quarter of heat predused is absorb by ice .

Grade:11

1 Answers

Arun
6 years ago
Dear Amar
 
E = mgh/4 = m Lf
h = 4 Lf/g
h = 4 *3.4 * 10^5/10
h = 13.6 * 104 metres
h = 136 km
 
Regards
Arun (askIITians forum expert)